A wild pig attacks a hunter with a velocity of 5 m/s. At the instant when the pig is at a distance of 100 m the hunter shoots an arrow with an angle of 40° above the horizontal.What is the initial velocity of the arrow to kill the wild pig?
Answer #1

5 m/s 670° 100m

let when arrow will hit the pig then distance from pig to hunter will be 100-x that is equal to range of arrow and let initial velocity of arrow is V

now X = Vpigt = 5*t ..............................(1)

here t is equal to time of flight of arrow

so now we apply projectile eqns

R=u2sin2\theta/g

t=2usin\theta/g

now 100-x=R

\Rightarrow100- Vpigt = R

using formula of R and t and g=9.8 m/s2

\Rightarrow100- 5*2Vsin40o/9.8 = V2sin80o/9.8

\Rightarrow100-0.656V=V20.1

\RightarrowV2+6.56V-1000=0

solving this eqn we get V= 28.51 m/sec